two angels are complementary. The sum measures of the first angel and half of the second angel is 64 degrees. find the measurments of angels.
Answer by ankor@dixie-net.com(15746)

two angels are complementary.
A + B = 90
The sum measures of the first angel and half of the second angel is 64 degrees.
A + .5B = 64
:
Subtract the 2nd eq from the 1st eq:
A + B = 90
A +.5B =64
------------subtraction eliminates A, find B
.5B = 26
B = 2*26
B = 52 degrees
then
A = 90 - 52
A = 38 degrees
:
:
Check this
38 + .5(52) = 64